# Adaptive and Efficient Learning with Blockwise Missing and Semi-Supervised Data

### v1

@@ 3.3 Relative Efficiency
${\rm AVar}(\hat{\beta}_{1})={\rm Var}\left[\boldsymbol{c}^{\top}\bar{\boldsymbol{S}}(\boldsymbol{X},Y)\big{|}\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y\right]+\frac{{\rm Var}\left\{{\rm E}\left[\boldsymbol{c}^{\top}\bar{\boldsymbol{S}}(\boldsymbol{X},Y)\big{|}\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y\right]\right\}}{1+{\bar{\rho}}}.$
@@
with a strict $>$ holding when the correlation of $\bar{S}_{1}(\boldsymbol{X},Y)$ and $\bar{\varphi}_{2}^{\prime}(\boldsymbol{X})$ is not zero. When both ${\mu}(\boldsymbol{X}_{\mathcal{P}_{m}}|\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y;\bar{\boldsymbol{\eta}})={\rm P}(\boldsymbol{X}_{\mathcal{P}_{m}}|\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y)$ and $\nu(Y|\boldsymbol{X};\bar{\boldsymbol{\theta}})={\rm P}(Y|\boldsymbol{X})$, we further have
${\rm AVar}(\hat{\beta}_{2})={\rm Var}\left\{\boldsymbol{c}^{\top}\bar{\boldsymbol{S}}(\boldsymbol{X},Y)-{\rm E}\left[\boldsymbol{c}^{\top}\bar{\boldsymbol{S}}(\boldsymbol{X},Y)\big{|}\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y\right]\big{|}\boldsymbol{X}\right\}+\frac{{\rm Var}\left\{{\rm E}\left[\boldsymbol{c}^{\top}\bar{\boldsymbol{S}}(\boldsymbol{X},Y)\big{|}\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y\right]\right\}}{1+{\bar{\rho}}}.$
In this case, $\mbox{RE}(\hat{\beta}_{2}|\hat{\beta}_{1})$ will monotonically increase with $\bar{\rho}$ and we will have ${\rm AVar}(\hat{\beta}_{2})<\sigma_{\sf ss}^{2}:={\rm Var}\left\{\boldsymbol{c}^{\top}\bar{\boldsymbol{S}}(\boldsymbol{X},Y)\big{|}\boldsymbol{X}\right\}$, where $\sigma_{\sf ss}^{2}$ is the semiparametric efficiency bound for the estimation of $\boldsymbol{c}^{\top}\bar{\boldsymbol{\gamma}}$ under the SS setting with only $\mathcal{L}\mathcal{C}$ and $\mathcal{U}\mathcal{C}$.
@@ 4.1 Settings and Benchmarks
$\boldsymbol{X}_{\mathcal{P}_{m}}=\begin{cases}\begin{pmatrix}0.3&-0.6&0.9\\ 0.6&-0.6&-0.3\end{pmatrix}(X_{1}^{2},X_{2}^{2},X_{3}^{2})^{\top}+\begin{pmatrix}-0.5\\ 0.5\end{pmatrix}+\epsilon&\mbox{ if }Y=1\\ \begin{pmatrix}-0.3&0.9&-0.6\\ -0.9&0.6&0.6\end{pmatrix}(X_{1}^{2},X_{2}^{2},X_{3}^{2})^{\top}+\epsilon&\mbox{ if }Y=0\\ \end{cases}$
@@
In the settings (I) and (II), the outcome model is the logistic model for $Y\sim\boldsymbol{X}$. We also include (III) and (IV) with linear models for $Y$ to facilitate the comparison with a broader set of existing methods for BM regression. In (I) and (III), the hyperparameters $\omega$ and $\rho$ are introduced and varied to study how the efficiency gain of our estimators changes according to the dependence between the missing and the observed variables as well as to the sample size ratio between $\mathcal{L}\mathcal{M}$ and $\mathcal{L}\mathcal{C}$. In (II), the hyperparameters $\omega$ and $\rho$ are fixed, and the dependence of missing conditioning on the observed variable is introduced through the quadratic function. In (IV), the hyperparameters $\omega$ and $\rho$ are fixed, and the more complex missing structure is brought into the simulation to appraise the robustness and efficiency of the estimators. In each scenario, we evaluate and compare the accuracy and efficiency of the estimators through their bias and mean square errors on each coefficient or averaged over all coefficients.
@@
We summarize all methods under comparison with some implementation details. $\mathcal{L}\mathcal{C}$-only: The standard estimator $\tilde{\boldsymbol{\gamma}}$ obtained using $\mathcal{L}\mathcal{C}$ data only. DEFUSE1: Our proposed estimator $\hat{\beta}_{1}$ based on $\mathcal{L}\mathcal{C}$ and $\mathcal{L}\mathcal{M}$ data. DEFUSE: Our final estimator $\hat{\beta}_{2}$. MICE: Multiple imputation by chained equations with the default settings and tuning parameters of the R package mice (van Buuren and Groothuis-Oudshoorn,, 2011). MBI: The multiple blockwise imputation method proposed by Xue and Qu, (2021) combining $\mathcal{L}\mathcal{C}$ and $\mathcal{L}\mathcal{M}$ data. HTLGMM: The data fusion approach proposed by Zhao et al., (2023) based on the generalized method of moments. SSB: The SS and BM fusion method proposed by Song et al., (2024) using the $\mathcal{L}\mathcal{C}$, $\mathcal{L}\mathcal{M}$ and $\mathcal{U}\mathcal{C}$ samples. Benchmarks MBI, HTLGMM, and SSB are all implemented under our low-dimensional regression settings without using any sparse penalties. SemiEff: The classic semiparametric efficient estimator $\hat{\beta}_{\sf semieff}$ constructed following Robins et al., (1994). SSL: The SS estimator using $\mathcal{L}\mathcal{C}$ and $\mathcal{U}\mathcal{C}$ (Gronsbell et al.,, 2022, e.g.), with its imputation model for $Y$ constructed through natural spline regression. We summarize all these methods in Table 1.
@@
Table 1: A list of the methods under comparison in our simulation studies. We also make a summary on their applicability to nonlinear (logistic) GLMs, whether they use the single or multiple BM data sets, and whether they use the unlabeled complete sample $\mathcal{U}\mathcal{C}$. All methods can be implemented and evaluated in our linear model and single BM setting (III) and some of them are only applicable to a subset of our simulation settings, e.g., MBI only applied to (III) and (IV) targeting the linear model.
For the conditional distribution of $\boldsymbol{X}_{\mathcal{P}_{m}}$ used in multiple approaches, including ours, we separately consider using both the parametric regression and the nonparametric kernel machine method to learn its conditional mean function and the standard method of moment to estimate its conditional covariance matrix as described in Remark 1. For ease of presentation, we will mainly report the results obtained using the parametric strategy and the results of kernel machine are provided in the Appendix. For parametric regression, we fit the ordinary multivariate linear regression for the missing covariates. In (I) and (II), we fit Gaussian linear models for $\boldsymbol{X}_{\mathcal{P}_{m}}\sim\boldsymbol{X}_{\mathcal{P}_{m}^{c}}$ stratified by the binary $Y$. In (III) and (IV), we fit Gaussian linear regression for $\boldsymbol{X}_{\mathcal{P}_{m}}\sim(\boldsymbol{X}_{\mathcal{P}_{m}^{c}},Y)$. In settings (I), (III), and (IV), our parametric model for $\boldsymbol{X}_{\mathcal{P}_{m}^{c}}$ is correctly specified while it is wrong under (II) due to the quadratic effects of $\boldsymbol{X}_{\mathcal{P}_{m}^{c}}$. We use the parametric modeling strategy for the conditional distribution of $Y$ used for imputation in our method. For the data-adaptive calibration functions, we set $w(\boldsymbol{x},y;\boldsymbol{\delta})=1+\boldsymbol{\delta}^{\top}(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ and $h(\boldsymbol{x},y;\boldsymbol{\zeta})=1+\zeta_{1}$ in all settings.
@@ 4.2 Comparison with existing methods
In Table 4, we present the RE (to $\mathcal{L}\mathcal{C}$-only) on each coefficient in the simulation setting (II) with a nonlinear model for $\boldsymbol{X}_{\mathcal{P}_{m}}$ not captured by the parametric conditional model. To demonstrate the importance of properly specifying the calibration function $w(\boldsymbol{x},y;\boldsymbol{\delta})$, we take $w(\boldsymbol{x},y;\boldsymbol{\delta})=1+\boldsymbol{\delta}^{\top}\xi(\boldsymbol{x},y)$ and include two choices on $\xi(\boldsymbol{x},y)$ for the construction of DEFUSE1: (1) $\xi(\boldsymbol{x},y)=1$ and (2) $\xi(\boldsymbol{x},y)=(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$. Choice $\xi(\boldsymbol{x},y)=1$ only involves an “intercept” term for calibration and is actually relevant to proposals in recent work like Miao et al., (2023). Interestingly, DEFUSE1 with $\xi(\boldsymbol{x},y)=(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ outperforms both SemiEff and DEFUSE1 with $\xi(\boldsymbol{x},y)=1$. For example, on $\gamma_{3}$, DEFUSE1 with $\xi(\boldsymbol{x},y)=(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ has around $50\%$ higher RE than both SemiEff and DEFUSE1 with $\xi(\boldsymbol{x},y)=1$. This demonstrates that our delicate data-adaptive calibration approach can effectively improve the estimation efficiency over the classic SemiEff when the imputation model $\tilde{\mu}_{\boldsymbol{X}_{\mathcal{P}_{m}}}$ is wrongly specified or of poor quality. To realize this advantage, it is important to specify the calibration function $w(\boldsymbol{x},y;\boldsymbol{\delta})$ in a proper way, which can be seen by comparing DEFUSE1 under the two different choices on $\xi(\boldsymbol{x},y)$ in this setting.
@@
DEFUSE1 $\xi(\boldsymbol{x},y)=1$ 3.48 3.24 2.62 0.99 0.98
DEFUSE1 $\xi(\boldsymbol{x},y)=(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ 4.35 4.35 3.59 1.21 1.11
SemiEff 3.53 3.13 2.35 0.98 0.99
SSL 1.00 1.10 0.99 0.93 1.38
HTLGMM 2.31 2.64 1.64 1.10 1.13
Table 4: Relative efficiency (RE) to $\mathcal{L}\mathcal{C}$-only on each coefficient in $\boldsymbol{\gamma}$ under Setting (II).
@@
Table 5: Relative efficiency (RE) to $\mathcal{L}\mathcal{C}$-only on each coefficient in $\boldsymbol{\gamma}$ under Setting (IV).
@@ 4.3 Patterns of relative efficiency
In Figures 3 and 5, we present the REs of DEFUSE1 to $\mathcal{L}\mathcal{C}$-only. A larger efficiency gain of DEFUSE1 over $\mathcal{L}\mathcal{C}$-only can be observed with an increasing sample size ratio $\rho=N_{\mathcal{L}\mathcal{M}}/N_{\mathcal{L}\mathcal{C}}$; see also Tables 2 and 3. This finding agrees with our Corollary 1 that $\mbox{RE}(\hat{\beta}_{1}|\boldsymbol{c}^{\top}\tilde{\boldsymbol{\gamma}})$ monotonically increases with $\rho$. On $\gamma_{5}$ corresponding to the partially missing $X_{5}$, the efficiency gain of DEFUSE1 tends to be higher with an increasing dependence parameter $\omega$. Intuitively, when $\omega=0$, $X_{5}$ will be independent from $(Y,\boldsymbol{X}_{\mathcal{P}_{m}^{c}})$ and the $\mathcal{L}\mathcal{M}$ sample with only $(Y,\boldsymbol{X}_{\mathcal{P}_{m}^{c}})$ observed cannot provide any information on $X_{5}$. Consequently, incorporating $\mathcal{L}\mathcal{M}$ could not help much on $\gamma_{5}$ whose score function is largely dependent on $X_{5}$. On the other hand, under a large $\omega$, $(Y,\boldsymbol{X}_{\mathcal{P}_{m}^{c}})$ observed on $\mathcal{L}\mathcal{M}$ can be informative to the score function of $\gamma_{5}$ and improve the estimation efficiency on it. Interestingly, for $\gamma_{1}$ and $\gamma_{3}$ corresponding to covariates in $\boldsymbol{X}_{\mathcal{P}_{m}^{c}}$ observed on $\mathcal{L}\mathcal{M}$, increasing $\omega$ tends to cause a lower efficiency gain of our method, which is opposite to the pattern of REs on $\gamma_{5}$. For example, in setting (III) with $\rho=1$, the RE of DEFUSE1 on $\gamma_{3}$ decrease from $4$ to $2$ when $\omega$ changes from $0.1$ to $0.8$; see Figure 5. This is because when $\omega$ becomes larger, the score functions of $\gamma_{1,3}$ tend to be more dependent on $\boldsymbol{X}_{\mathcal{P}_{m}}$, which makes the missingness of $\boldsymbol{X}_{\mathcal{P}_{m}}$ on $\mathcal{L}\mathcal{M}$ more impactful.

### v2

@@ 3.3 Relative Efficiency
${\rm AVar}(\hat{\sqrt{n}\beta}_{1})={\rm E}\left({\rm Var}\left[\boldsymbol{c}^{\top}\bar{\boldsymbol{S}}(\boldsymbol{X},Y)\big{|}\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y\right]\right)+\frac{{\rm Var}\left\{{\rm E}\left[\boldsymbol{c}^{\top}\bar{\boldsymbol{S}}(\boldsymbol{X},Y)\big{|}\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y\right]\right\}}{1+{\bar{\rho}}}.$
@@
with a strict $>$ holding when the correlation of $\bar{S}_{1}(\boldsymbol{X},Y)$ and $\bar{\varphi}_{2}^{\prime}(\boldsymbol{X})$ is not zero and $\mbox{RE}(\hat{\beta}_{2}|\hat{\beta}_{1})$ monotonically increases with $\bar{\rho}=\lim_{n\rightarrow\infty}\rho$. When ${\mu}(\boldsymbol{X}_{\mathcal{P}_{m}}|\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y;\bar{\boldsymbol{\eta}})={\rm P}(\boldsymbol{X}_{\mathcal{P}_{m}}|\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y)$ and $\nu(Y|\boldsymbol{X};\bar{\boldsymbol{\theta}})={\rm P}(Y|\boldsymbol{X})$, we further have that
${\rm AVar}(\sqrt{n}\hat{\beta}_{2})=$ ${\rm E}\left\{{\rm Var}\left[\boldsymbol{c}^{\top}\bar{\boldsymbol{S}}(\boldsymbol{X},Y)-\frac{\bar{\rho}}{1+\bar{\rho}}{\rm E}\left[\boldsymbol{c}^{\top}\bar{\boldsymbol{S}}(\boldsymbol{X},Y)\big{|}\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y\right]\big{|}\boldsymbol{X}\right]\right\}$
$+\frac{\bar{\rho}{\rm Var}\left\{{\rm E}\left[\boldsymbol{c}^{\top}\bar{\boldsymbol{S}}(\boldsymbol{X},Y)\big{|}\boldsymbol{X}_{\mathcal{P}_{m^{c}}},Y\right]\right\}}{(1+{\bar{\rho}})^{2}}$
${\rm AVar}(\sqrt{n}\hat{\beta}_{2})<\sigma_{\sf ss}^{2}:={\rm Var}\left\{\boldsymbol{c}^{\top}\bar{\boldsymbol{S}}(\boldsymbol{X},Y)\big{|}\boldsymbol{X}\right\}$, where $\sigma_{\sf ss}^{2}$ is the semiparametric efficiency bound for the estimation of $\boldsymbol{c}^{\top}\bar{\boldsymbol{\gamma}}$ under the SS setting with only $\mathcal{L}\mathcal{C}$ and $\mathcal{U}\mathcal{C}$.
@@ 4.1 Settings and Benchmarks
$\boldsymbol{X}_{\mathcal{P}_{m}}=\begin{cases}\begin{pmatrix}1&-2&3\\ 2&-2&-1\end{pmatrix}(X_{1}^{2},X_{2}^{2},X_{3}^{2})^{\top}+\begin{pmatrix}-1\\ 1\end{pmatrix}+\epsilon&\mbox{ if }Y=1\\ \begin{pmatrix}-1&3&-2\\ -3&2&2\end{pmatrix}(X_{1}^{2},X_{2}^{2},X_{3}^{2})^{\top}+\epsilon&\mbox{ if }Y=0\\ \end{cases}$
@@
In the settings (I) and (II), the outcome model is the logistic model for $Y\sim\boldsymbol{X}$. In (II), the dependence of $\boldsymbol{X}_{\mathcal{P}_{m}}$ on $\boldsymbol{X}_{\mathcal{P}_{m}^{c}}$ is quadratic, which may make the imputation models subject to model misspecification. We also include (III) and (IV) with linear models for $Y$ to facilitate the comparison with a broader set of existing methods for BM regression. In (I) and (III), the hyperparameters $\omega$ and $\rho$ are introduced and varied to study how the efficiency gain of our estimators changes according to the dependence between the missing and the observed variables as well as to the sample size ratio between $\mathcal{L}\mathcal{M}$ and $\mathcal{L}\mathcal{C}$. In (IV), the hyperparameters $\omega$ and $\rho$ are fixed, and the more complicated multiple BM structure is brought in. In each scenario, we evaluate and compare the estimators through their bias and mean square errors on each coefficient or averaged over all of them.
@@
We summarize all methods under comparison with some implementation details; see also Table 1. $\mathcal{L}\mathcal{C}$-only: The standard estimator $\tilde{\boldsymbol{\gamma}}$ obtained using $\mathcal{L}\mathcal{C}$ data only. DEFUSE1: Our proposed estimator $\hat{\beta}_{1}$ based on $\mathcal{L}\mathcal{C}$ and $\mathcal{L}\mathcal{M}$ data. DEFUSE: Our final estimator $\hat{\beta}_{2}$. MICE: Multiple imputation by chained equations with R package mice (van Buuren and Groothuis-Oudshoorn,, 2011). MBI: The multiple blockwise imputation method proposed by Xue and Qu, (2021) combining $\mathcal{L}\mathcal{C}$ and $\mathcal{L}\mathcal{M}$ data. HTLGMM: The data fusion approach proposed by Zhao et al., (2023) based on the generalized method of moments. SSB: The SS and BM fusion method proposed by Song et al., (2024) using the $\mathcal{L}\mathcal{C}$, $\mathcal{L}\mathcal{M}$ and $\mathcal{U}\mathcal{C}$ samples. Benchmarks MBI, HTLGMM, and SSB are all implemented under our low-dimensional regression settings without using any sparse penalties. SemiEff: The classic semiparametric efficient estimator $\hat{\beta}_{\sf semieff}$ constructed following Robins et al., (1994). SSL: The SS estimator using $\mathcal{L}\mathcal{C}$ and $\mathcal{U}\mathcal{C}$ (Gronsbell et al.,, 2022, e.g.), with its imputation model for $Y$ constructed through natural spline regression.
@@
Table 1: Methods under our comparison, including a summary on their applicability to nonlinear (logistic) GLMs, whether they use the single or multiple BM data sets, and whether they use the unlabeled complete sample $\mathcal{U}\mathcal{C}$. All methods can be implemented and evaluated in our linear model and single BM setting (III) and some are only applicable to a subset of our simulation settings, e.g., MBI only applied to (III) and (IV) with linear outcome models.
For the conditional distribution of $\boldsymbol{X}_{\mathcal{P}_{m}}$ used in multiple approaches, we carry out parametric regression to learn its conditional mean function and the standard method of moment to estimate its conditional covariance matrix as described in Remark 1. In (I) and (II), we fit Gaussian linear models for $\boldsymbol{X}_{\mathcal{P}_{m}}\sim\boldsymbol{X}_{\mathcal{P}_{m}^{c}}$ stratified by the binary $Y$. In (III) and (IV), we fit Gaussian linear regression for $\boldsymbol{X}_{\mathcal{P}_{m}}\sim(\boldsymbol{X}_{\mathcal{P}_{m}^{c}},Y)$. In settings (I), (III), and (IV), our parametric model for $\boldsymbol{X}_{\mathcal{P}_{m}^{c}}$ is correctly specified while it is wrong under (II) due to the quadratic effects of $\boldsymbol{X}_{\mathcal{P}_{m}^{c}}$. We also fit parametric models for imputation of $Y$. For the data-adaptive calibration functions used in DEFUSE, we set $w(\boldsymbol{x},y;\boldsymbol{\delta})=1+\boldsymbol{\delta}^{\top}(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ and $h(\boldsymbol{x},y;\boldsymbol{\zeta})=1+\zeta_{1}$ in all settings.
@@ 4.2 Comparison with existing methods
In Tables 3 and 5, we present the RE (to $\mathcal{L}\mathcal{C}$-only) on each coefficient in the simulation setting (II) with a nonlinear model for $\boldsymbol{X}_{\mathcal{P}_{m}}$ not captured by the parametric conditional model. To demonstrate the importance of properly specifying the calibration function $w(\boldsymbol{x},y;\boldsymbol{\delta})$, we take $w(\boldsymbol{x},y;\boldsymbol{\delta})=1+\boldsymbol{\delta}^{\top}\xi(\boldsymbol{x},y)$ and include two choices on $\xi(\boldsymbol{x},y)$ for the construction of DEFUSE1: (1) $\xi(\boldsymbol{x},y)=1$ and (2) $\xi(\boldsymbol{x},y)=(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$. Choice $\xi(\boldsymbol{x},y)=1$ only involves an “intercept” term for calibration and is actually relevant to proposals in recent work like Miao et al., (2023). Interestingly, DEFUSE1 with $\xi(\boldsymbol{x},y)=(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ outperforms both SemiEff and DEFUSE1 with $\xi(\boldsymbol{x},y)=1$. For example, on $\gamma_{3}$, DEFUSE1 with $\xi(\boldsymbol{x},y)=(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ has around $50\%$ higher RE than both SemiEff and DEFUSE1 with $\xi(\boldsymbol{x},y)=1$. This demonstrates that our new data-adaptive calibration approach can effectively improve the estimation efficiency over the classic SemiEff when the imputation model $\tilde{\mu}_{\boldsymbol{X}_{\mathcal{P}_{m}}}$ is wrongly specified or of poor quality. To realize this advantage, it is important to specify the calibration function $w(\boldsymbol{x},y;\boldsymbol{\delta})$ in a proper way, which can be seen by comparing DEFUSE1 under the two different choices on $\xi(\boldsymbol{x},y)$ in this setting.
@@
Methods ratio $\rho$ = 2 ratio $\rho$ = 3 ratio $\rho$ = 4 ratio $\rho$ = 5
RE Bias RE Bias RE Bias RE Bias
SemiEff 1.88 0.020 2.22 0.010 2.52 0.016 2.63 0.019
DEFUSE1 $\xi(\boldsymbol{x},y)=1$ 1.91 0.018 2.26 0.018 2.57 0.024 2.70 0.022
DEFUSE1 $\xi(\boldsymbol{x},y)=(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ 2.09 0.013 2.51 0.021 2.98 0.022 3.23 0.027
MICE 2.22 0.103 2.72 0.135 3.66 0.151 4.11 0.167
SSL 1.04 0.47 1.06 0.47 1.04 0.45 1.05 0.046
HTLGMM 1.44 0.038 1.63 0.037 1.78 0.042 1.85 0.046
Table 3: Average absolute bias and relative efficiency (RE) to $\mathcal{L}\mathcal{C}$-only of the methods in simulation setting (II) with $\omega=1$ and $\rho\in\{2,3,4,5\}$.
@@
SemiEff 4.48 3.84 2.86 0.99 1.00
DEFUSE1 $\xi(\boldsymbol{x},y)=1$ 4.44 3.99 3.12 0.98 0.96
DEFUSE1 $\xi(\boldsymbol{x},y)=(1,\boldsymbol{x}_{\mathcal{P}_{m}^{c}},y)$ 4.87 5.11 3.90 1.23 1.06
MICE 4.86 10.01 4.92 0.18 0.58
SSL 1.01 1.08 1.00 0.93 1.25
HTLGMM 2.41 2.80 1.90 1.06 1.09
Table 5: Current Relative efficiency (RE) to $\mathcal{L}\mathcal{C}$-only on each coefficient under Setting (II).
@@
Table 6: Relative efficiency (RE) to $\mathcal{L}\mathcal{C}$-only on each coefficient under Setting (IV).
@@ 4.3 Patterns of relative efficiency
In Figures 3 and 5, we present the REs of DEFUSE1 to $\mathcal{L}\mathcal{C}$-only. A larger efficiency gain of DEFUSE1 over $\mathcal{L}\mathcal{C}$-only can be observed with an increasing sample size ratio $\rho=N_{\mathcal{L}\mathcal{M}}/N_{\mathcal{L}\mathcal{C}}$; see also Tables 2 and 4. This finding agrees with our Corollary 1 that $\mbox{RE}(\hat{\beta}_{1}|\boldsymbol{c}^{\top}\tilde{\boldsymbol{\gamma}})$ monotonically increases with $\rho$. On $\gamma_{5}$ corresponding to the partially missing $X_{5}$, the efficiency gain of DEFUSE1 tends to be higher with an increasing dependence parameter $\omega$. Intuitively, when $\omega=0$, $X_{5}$ will be independent from $(Y,\boldsymbol{X}_{\mathcal{P}_{m}^{c}})$ and the $\mathcal{L}\mathcal{M}$ sample with only $(Y,\boldsymbol{X}_{\mathcal{P}_{m}^{c}})$ observed cannot provide any information on $X_{5}$. Consequently, incorporating $\mathcal{L}\mathcal{M}$ could not help much on $\gamma_{5}$ whose score function is largely dependent on $X_{5}$. On the other hand, under a large $\omega$, $(Y,\boldsymbol{X}_{\mathcal{P}_{m}^{c}})$ observed on $\mathcal{L}\mathcal{M}$ can be informative to the score function of $\gamma_{5}$ and improve the estimation efficiency on it. Interestingly, for $\gamma_{1}$ and $\gamma_{3}$ corresponding to covariates in $\boldsymbol{X}_{\mathcal{P}_{m}^{c}}$ observed on $\mathcal{L}\mathcal{M}$, increasing $\omega$ tends to cause a lower efficiency gain of our method, which is opposite to the pattern of REs on $\gamma_{5}$. For example, in setting (III) with $\rho=1$, the RE of DEFUSE1 on $\gamma_{3}$ decrease from $4$ to $2$ when $\omega$ changes from $0.1$ to $0.8$; see Figure 5. This is because when $\omega$ becomes larger, the score functions of $\gamma_{1,3}$ tend to be more dependent on $\boldsymbol{X}_{\mathcal{P}_{m}}$, which makes the missingness of $\boldsymbol{X}_{\mathcal{P}_{m}}$ on $\mathcal{L}\mathcal{M}$ more impactful.